OpenStudy (anonymous):

Integrate ∫(2x4−5)6x3dx.

geerky42 (geerky42):

Like this? \(\Large \displaystyle \int\limits \left( 2x^4 - 5 \right)6x^3 \space \text{d}x\)

OpenStudy (anonymous):

it's not like that

geerky42 (geerky42):

Then what is it?

geerky42 (geerky42):

use ^ for power.